# Reliable quantum advantage in quantum battery charging

### v1

@@ I Introduction
A quantum battery is a generic quantum system capable of storing energy after being charged, and then delivering it on demand when needed [15, 14]. A vast theoretical literature has become available on this topic, mostly concerning theoretical proposals [16, 17, 18, 19, 20, 21, 22, 23], followed only recently by the first experimental implementations [24, 25, 26, 27]. Nevertheless, the applications and the role of quantum batteries in future technology with the aim, e.g., of enhancing the capacity of energy storage devices [28], or becoming an integrated part of quantum computing processes [29], still need to be thoroughly investigated. On the other hand, the interest in quantum batteries is also motivated by the usefulness that such theoretical models represent in fundamental physics studies, e.g., in understanding the role of energy fluctuations in quantum systems by using thermodynamic-based approaches.
@@
Before proceeding, let us notice that models of a JC quantum battery have already been addressed in the literature. Among the different characterizations, an analysis similar to ours has been reported in Ref. [21], although with the main purpose of charging an oscillator (a micromaser quantum battery) via a stream of qubits, each one prepared in their excited state. Another recent work has analyzed the charge of multiple three-level systems coupled to a cavity mode prepared in Fock, coherent, or squeezed coherent states, respectively [22], obtaining conclusions that are somehow opposite to ours, in the sense that the coherent state would appear to be the one that brings some advantage compared to the others. Finally, sticking to the single-qubit JC model, a two-photon JC quantum battery was investigated in Ref. [23], where it has been shown that the use of a Fock state allows to inject a larger amount of energy as compared to the classical counterparts (coherent and squeezed states), in line with the main results obtained in the present work.
@@ II The Model
We consider the case of a qubit with internal transition frequency $\omega_{\text{qub}}$, interacting for a finite time $\tau$ with a single mode of an electromagnetic field confined to a resonator, whose harmonic oscillating frequency is denoted as $\omega_{\text{cav}}$. The interaction between the qubit and the cavity mode occurs, in general, through a time-dependent coupling strength $g(t)$, which in its simplest version can be set to be nonzero (and constant within the time of flight into the resonator) only between time $t=0$ and time $t=\tau$, i.e., explicitly we have $g(t)=g\cdot\Theta(t)\Theta(\tau-t)$. With this approximation in mind, therefore, we can think of the coupling strength as a constant parameter for times $0\leq t\leq\tau$, which leads to a theoretical description in terms of the Rabi Hamiltonian [33], in general. However, if we restrict the analysis to the regime in which $g/\omega_{0}\ll 1$, with $\omega_{\text{qub}}\simeq\omega_{0}\simeq\omega_{\text{cav}}$, then the counter-rotating terms in the Rabi Hamiltonian can be neglected (the so-called rotating wave approximation), leaving us with the simple JC Hamiltonian [34, 35, 36, 37, 38, 39]:
@@ III Dynamical fluctuations and Full Counting Statistics
As an example, if we identify $\hat{A}=\hat{H}_{\text{qub}}$ (i.e., the qubit Hamiltonian), then $\Delta a_{\tau}\rightarrow\Delta U_{\tau}$ is a stochastic variable that represents the qubit energy variation $\Delta U$ over the time window $\tau$, associated with two sequential energy measurements performed on the qubit at different times. As the notations suggests, we will denote the qubit energy variation with $\Delta U_{\tau}$, that is the internal energy variation of the system of interest (i.e., the battery). Besides, we will consider as heat the change in energy of the environment [45]: in this case, it is the energy change of the cavity (i.e., the charger), which actually acts as a sort of environment. Therefore, by identifying $\hat{A}$ with $\hat{H}_{\text{cav}}$, the heat is $\Delta a_{\tau}\rightarrow Q_{\tau}$, which is linked to two sequential energy measurements made on the cavity. Finally, by employing the first law of thermodynamics, we get the work $W_{\tau}$ as the energy contribution stemming from the energy conservation, i.e., $\Delta U_{\tau}=Q_{\tau}+W_{\tau}$. This contribution can be interpreted as the amount of energy stored in the interaction between the two systems: as it will be shown in Eqs. (12), $W_{\tau}$ vanishes in case of perfect resonance (i.e., whenever the JC model is energy conserving), while it is nonzero if $\Delta\omega\neq 0$.
@@
Once the generating function is known, it is possible to find the statistical moments of $\Delta U_{\tau}$ and $Q_{\tau}$, by performing the derivative of $\mathscr{G}_{\tau}(\boldsymbol{\chi})$ with respect to $\chi_{1}$ and $\chi_{1}$, respectively, and by setting $\chi_{1}=0=\chi_{2}$. The statistics of the work $W_{\tau}$ can then be obtained by using the first law of thermodynamics:
@@ IV.1 The signal-to-noise ratio
Figure 2: Plot of the $\text{SNR}(\Delta U_{\tau})$ from Eq. (6) as a function of the time window $g\tau$, related to four different initial cavity states: pure Fock (continuous green line), coherent (dashed red), thermal (dotted blue), and squeezed coherent (dash-dotted orange). Importantly for a fair comparison between them, all these states have the same average photon number, which is $\langle n\rangle=N=5$. Moreover, here we set the other parameters to be $\omega_{\text{qub}}=1\text{ GHz}$ $g/\omega_{\text{qub}}=10^{-2}$, $\Delta\omega=\omega_{\text{qub}}-\omega_{\text{cav}}$ and $\Delta\omega/\omega_{\text{qub}}=5\cdot 10^{-3}$. As reported in Eq. (13), $\text{SNR}(\Delta U_{\tau})$, $\text{SNR}(Q_{\tau})$, and $\text{SNR}(W_{\tau})$ are quantitatively identical.
@@
Analogously, we can calculate the SNR for different initial cavity states, e.g., the ones defined in Eq. (2). Although an analytical formulation is still possible, in these cases it involves summations over infinite terms (i.e., over the cavity energy levels): therefore, computations must be carried out numerically, by truncating the summations. The results corresponding to four different initial cavity mode states are reported in Fig. 2. Here we have fixed $N=5$, and consequently we choose $\beta=\log(\frac{N+1}{N})/\hbar\omega_{\text{cav}}$ (by direct application of the Bose-Einstein statistical relation), and $\alpha=\sqrt{N}$; regarding the squeezed coherent state, we choose a real squeezing parameter $\zeta=0.6$, and we picked $\tilde{\alpha}\simeq 3.906$. This way, each Gaussian state contains $\langle n\rangle=N=5$ photons on average. This way, a direct comparison can be made between the different cases, showing that an effective quantum advantage in the charging process of a JC quantum battery is achieved when the cavity mode is prepared in a Fock state as compared to any Gaussian (i.e., thermal, coherent, or squeezed coherent) states of light, as anticipated at the beginning of this Section.
@@ IV.2 Charging protocol: power and charge efficacy
The simulated results showing the produced power and the fidelity evaluated over a time window $g\tau$ are depicted in Fig. 3. First of all, we immediately notice how the Fock state cavity produces slightly more power than the three Gaussian states. (b) Fidelity $F$, as defined in Eq. (17), between $\hat{\rho}_{\text{qub}}(\tau)$ and $\hat{\rho}_{\text{qub}}^{\text{exc}}$, where $\hat{\rho}_{\text{qub}}^{\text{exc}}=\ket{e}\bra{e}$ |. The optimal charge (i.e., the process for which the final qubit state approaches |$\ket{e}\bra{e}$ |), can be obtained only when the cavity is prepared in a Fock state. In that case, $F\to 1$ when the qubit and the cavity mode are perfectly resonant, and when the charging time is the optimal time of Eq. (15). It is worth emphasising that the charging efficacy is maximized at the very same value of time $\tau_{\text{opt}}$ that optimized the SNR. This fact, far from obvious, strongly stresses the importance of incorporating the role of energy fluctuations into the picture when characterizing the performance of thermodynamic processes.
@@ VI Conclusions
As a final consideration, we reckon that these findings can be particularly suitable for the thermodynamic analysis of a wide variety of quantum processes, especially the ones for which energy fluctuations play a prominent role. Moreover, they provide a contribution to recent results concerning the investigation of the energetics of bipartite systems [55, 56], which fit well with the model discussed in the present work.
@@ Acknowledgements
The authors are grateful to S. P. Prasad, M. Maffei, A. Auffèves, M. Kolář, L. Slodička, G. Benenti, and L. Razzoli for the fruitful scientific discussions. DR acknowledges the kind hospitality of the Majulab at the Centre for Quantum Technologies in Singapore, as well as the warm welcome of the Optics Department at the Palacký University in Olomouc. DG acknowledges the ‘National Quantum Science Technology Institute’ (NQSTI, PE4, Spoke 1) within the PNRR project PE0000023. GG acknowledges funding from the Italian Ministry of Research (MUR) under the “Rita Levi Montalcini” project THESEUS. RF acknowledges the grant 21-13491X from the Czech Science Foundation.
@@ Appendix A The JC evolution operator
$\begin{split}\hat{\mathcal{U}}_{00}&=e^{-\frac{i}{2}\omega_{\text{cav}}(\hat{\mathds{I}}+2\hat{a}^{\dagger}\hat{a})\tau}\biggl{[}\mathcal{C}(\tau,\hat{\varphi}+g^{2}\hat{\mathds{I}})-i\frac{\Delta\omega}{2}\mathcal{S}(\tau,\hat{\varphi}+g^{2}\hat{\mathds{I}})\biggr{]}\\ \hat{\mathcal{U}}_{01}&=-ige^{-\frac{i}{2}\omega_{\text{cav}}\tau}e^{-i\omega_{\text{cav}}\hat{a}^{\dagger}\hat{a}\tau}\hat{a}\mathcal{S}(\tau,\hat{\varphi})\\ \hat{\mathcal{U}}_{10}&=-ige^{+\frac{i}{2}\omega_{\text{cav}}\tau}e^{-i\omega_{\text{cav}}\hat{a}^{\dagger}\hat{a}\tau}\mathcal{S}(\tau,\hat{\varphi})\hat{a}^{\dagger}\\ \hat{\mathcal{U}}_{11}&=e^{+\frac{i}{2}\omega_{\text{cav}}\tau}e^{-i\omega_{\text{cav}}\hat{a}^{\dagger}\hat{a}\tau}\biggl{[}\mathcal{C}(\tau,\hat{\varphi})+i\frac{\Delta\omega}{2}\mathcal{S}(\tau,\hat{\varphi})\biggr{]}\end{split}$ (18)
@@ Appendix B The generating function of the statistics
$\begin{split}&\mathcal{G}_{\tau}(\boldsymbol{\chi})=\rho_{00}(0)\langle\zeta(\hat{a}^{\dagger}\hat{a}+\hat{\mathds{I}})\rangle_{\text{cav}}\\ &+ig\rho_{01}(0)e^{+i\frac{\chi_{1}}{2}\hbar\omega_{\text{qub}}+i\frac{\chi_{2}}{2}\hbar\omega_{\text{cav}}}\langle\eta(\hat{a}^{\dagger}\hat{a})\hat{a}^{\dagger}\rangle_{\text{cav}}\\ &-ig\rho_{10}(0)e^{+i\frac{\chi_{1}}{2}\hbar\omega_{\text{qub}}+i\frac{\chi_{2}}{2}\hbar\omega_{\text{cav}}}\langle\hat{a}\eta^{\ast}(\hat{a}^{\dagger}\hat{a})\rangle_{\text{cav}}\\ &+g^{2}\rho_{11}(0)e^{+i\chi_{1}\hbar\omega_{\text{qub}}+i\chi_{2}\hbar\omega_{\text{cav}}}\langle\mathscr{S}^{2}(\tau,\hat{a}^{\dagger}\hat{a})\hat{a}^{\dagger}\hat{a}\rangle_{\text{cav}}\\ &+g^{2}\rho_{00}(0)e^{-i\chi_{1}\hbar\omega_{\text{qub}}-i\chi_{2}\hbar\omega_{\text{cav}}}\langle\hat{a}\mathscr{S}^{2}(\tau,\hat{a}^{\dagger}\hat{a})\hat{a}^{\dagger}\rangle_{\text{cav}}\\ &-ig\rho_{01}(0)e^{-i\frac{\chi_{1}}{2}\hbar\omega_{\text{qub}}-i\frac{\chi_{2}}{2}\hbar\omega_{\text{cav}}}\langle\eta(\hat{a}^{\dagger}\hat{a})\hat{a}^{\dagger}\rangle_{\text{cav}}\\ &+ig\rho_{10}(0)e^{-i\frac{\chi_{1}}{2}\hbar\omega_{\text{qub}}-i\frac{\chi_{2}}{2}\hbar\omega_{\text{cav}}}\langle\hat{a}\eta^{\ast}(\hat{a}^{\dagger}\hat{a})\rangle_{\text{cav}}\\ &+\rho_{11}(0)\langle\zeta(\hat{a}^{\dagger}\hat{a})\rangle_{\text{cav}}\end{split}$ (23)
@@
and $\langle\hat{A}\rangle_{\text{cav}}\equiv\text{Tr}_{\text{cav}}\bigl{\{}\hat{A}(\tau)\hat{\rho}_{\text{cav}}(0)\bigr{\}}$.
Expression (23) greatly simplifies if we choose, for example, $\hat{\rho}_{\text{qub}}(0)=\ket{g}\bra{g}$ |: by adopting the convention |$\ket{e}=\footnotesize\begin{pmatrix}1\\ 0\end{pmatrix}$ and |$\ket{g}=\footnotesize\begin{pmatrix}0\\ 1\end{pmatrix}$, indeed, $\rho_{ij}(0)=1\iff i=j=1$, otherwise $\rho_{ij}(0)=0$ for $i\neq j$ or $i=j=0$. This is actually the case of the charging protocol of the JC quantum battery, where the qubit starts in the ground state. Now, let we further assume that the initial state of the cavity is a Fock state $\hat{\rho}_{\text{cav}}(0)=\ket{N}\bra{N}$ | with $N$ photons, and let us set $\chi_{2}=0$ (for example, because we may be interested in computing the statistics of the only $\Delta U_{\tau}$). In that case, we get Eq. (7). This can be easily seen from Eq. (23), which becomes
$\begin{split}\mathcal{G}_{\tau}(\boldsymbol{\chi})&=g^{2}e^{+i\chi_{1}\hbar\omega_{\text{qub}}}\langle\mathscr{S}^{2}(\tau,\hat{a}^{\dagger}\hat{a})\hat{a}^{\dagger}\hat{a}\rangle_{\text{cav}}\\ &\quad+\langle\zeta(\hat{a}^{\dagger}\hat{a})\rangle_{\text{cav}}\end{split}$ (25)
because the only non-vanishing factor is $\rho_{11}(0)$. The operator-valued expression $\zeta(\hat{a}^{\dagger}\hat{a})$, defined in Eq. (24), as well as the expression $\mathscr{S}^{2}(\tau,\hat{a}^{\dagger}\hat{a})\hat{a}^{\dagger}\hat{a}$, are then traced over the cavity’s degrees of freedom, together with the initial state of the cavity. For what concerns $\zeta(\hat{a}^{\dagger}\hat{a})$, we have $\langle\zeta(\hat{a}^{\dagger}\hat{a})\rangle_{\text{cav}}=\text{Tr}\{\zeta(\hat{a}^{\dagger}\hat{a})\ket{N}\bra{N}\}$ | }, that translates into $\bra{N}\zeta(\hat{a}^{\dagger}\hat{a})\ket{N}$. It is then straightforward to evaluate $\bra{N}\mathcal{S}^{2}(\tau,\hat{\varphi})\ket{N}=\mathscr{S}^{2}(\tau)$ and $\bra{N}\mathcal{C}^{2}(\tau,\hat{\varphi})\ket{N}=\mathscr{C}^{2}(\tau)$, where the functions $\mathscr{S}(\tau)$ and $\mathscr{C}(\tau)$ are respectively defined as:

### v2

@@ I Introduction
A quantum battery is a generic quantum system capable of storing energy after being charged, and then delivering it on demand when needed [15, 14]. A vast theoretical literature has become available on this topic, mostly concerning theoretical proposals [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26], followed only recently by the first experimental implementations [27, 28, 29, 30]. Nevertheless, the applications and the role of quantum batteries in future technology with the aim, e.g., of enhancing the capacity of energy storage devices [31], or becoming an integrated part of quantum computing processes [32], still need to be thoroughly investigated. On the other hand, the interest in quantum batteries is also motivated by the usefulness that such theoretical models represent in fundamental physics studies, e.g., in understanding the role of energy fluctuations in quantum systems by using thermodynamic-based approaches. While different routes can be taken (such as the investigation of upper bounds on non-equilibrium fluctuation of observables in quantum dynamics [33, 34]), quantum batteries are a promising test field for quantum thermodynamics itself.
@@
Before proceeding, let us notice that models of a JC quantum battery have already been addressed in the literature. Among the different characterizations, an analysis similar to ours has been reported in Ref. [21], although with the main purpose of charging an oscillator (a micromaser quantum battery) via a stream of qubits, each one prepared in their excited state. Similarly, three-level systems in place of qubits were employed in Ref. [25], where the possibility of swapping the role of battery and charger was also investigated. Sticking to the single-qubit JC model, a two-photon JC quantum battery was investigated in Ref. [23], where it has been shown that the use f a Fock state allows to inject a larger amount of energy as compared to the classical counterparts (coherent and squeezed states), in line with the main results obtained in the present work. We also emphasize that the advantage of using a Fock state might depend on the physical situation that has been taken into account. As an example, another recent work analyzed the charge of multiple three-level systems coupled to a cavity mode prepared in Fock, coherent, or squeezed coherent states, respectively [22]. In the latter results, the coherent state brings some advantage when compared to the others. However, the advantage provided by the coherent state is likely due to collective effects, such as the unwanted generation of entanglement between the systems, which seems to increase when using Fock or squeezed states, thus preventing the possibility for these states to enhance the battery performances. On the contrary, our model focuses on the interaction between the cavity and a single qubit at a time, which does not allow the emergence of quantum correlations between subsequent qubits flying into the cavity region. In such a context, Fock states seem to bring a quantitative advantage over the others.
@@ II The Model
We consider the case of a qubit with internal transition frequency $\omega_{\text{qub}}$, interacting for a finite time $\tau$ with a single mode of an electromagnetic field confined to a resonator, whose harmonic oscillating frequency is denoted as $\omega_{\text{cav}}$. The interaction between the qubit and the cavity mode occurs, in general, through a time-dependent coupling strength $g(t)$, which in its simplest version can be set to be nonzero (and constant within the time of flight into the resonator) only between time $t=0$ and time $t=\tau$, i.e., explicitly we have $g(t)=g\cdot\Theta(t)\Theta(\tau-t)$, where $\Theta(t)$ is the Heaviside step function: $\Theta(t)=1$ for $t\geq 0$ and $\Theta(t)=0$ elsewhere. With this approximation in mind, therefore, we can think of the coupling strength as a constant parameter for times $0\leq t\leq\tau$, which leads to a theoretical description in terms of the Rabi Hamiltonian [38], in general. However, if we restrict the analysis to the regime in which $g/\omega_{0}\ll 1$, with $\omega_{\text{qub}}\simeq\omega_{0}\simeq\omega_{\text{cav}}$, then the counter-rotating terms in the Rabi Hamiltonian can be neglected (the so-called rotating wave approximation [39]), leaving us with the simple JC Hamiltonian [39, 40, 41, 42, 43, 44, 45]:
@@ III Dynamical fluctuations and Full Counting Statistics
As an example, if we identify $\hat{A}=\hat{H}_{\text{qub}}$ (i.e., the qubit Hamiltonian), then $\Delta a_{\tau}\rightarrow\Delta U_{\tau}$ is a stochastic variable that represents the qubit energy variation $\Delta U$ over the time window $\tau$, associated with two sequential energy measurements performed on the qubit at different times. As the notations suggests, we will denote the qubit energy variation with $\Delta U_{\tau}$, that is the internal energy variation of the system of interest (i.e., the battery). Besides, we will consider as heat the change in energy of the environment [50]: in this case, it is the energy change of the cavity (i.e., the charger), which actually acts as a sort of environment. Therefore, by identifying $\hat{A}$ with $\hat{H}_{\text{cav}}$, the heat is $\Delta a_{\tau}\rightarrow Q_{\tau}$, which is linked to two sequential energy measurements made on the cavity. Finally, by employing the first law of thermodynamics, we get the work $W_{\tau}$ as the energy contribution stemming from the energy conservation, i.e., $\Delta U_{\tau}=Q_{\tau}+W_{\tau}$. This contribution can be interpreted as the amount of energy stored in the correlations between the two systems, in analogy with the expectation value of the interaction Hamiltonian at time $t=\tau$, $\langle\hat{H}_{\text{int}}(\tau)\rangle$. Anyhow, the latter is no more than an analogy: as discussed before, $\langle W_{\tau}\rangle$ is the average value of a process-dependent stochastic variable over a time window $\tau$, while $\langle\hat{H}_{\text{int}}(\tau)\rangle$ is a state-dependent quantum mechanical expectation value at time $t=\tau$. Although they are quantitatively identical at the level of average values, their interpretation is conceptually different, indeed. We notice, for example, that the variance of $W_{\tau}$ does not coincide with the state-dependent variance $\langle\hat{H}^{2}_{\text{int}}(\tau)\rangle-\langle\hat{H}_{\text{int}}(\tau)\rangle^{2}$, in general.
As it will be shown in Eqs. (12), $W_{\tau}$ vanishes in case of perfect resonance (i.e., whenever the JC model is energy conserving), while it is nonzero if $\Delta\omega\neq 0$.
@@
Once the generating function is known, it is possible to find the statistical moments of $\Delta U_{\tau}$ and $Q_{\tau}$, by performing the derivative of $\mathscr{G}_{\tau}(\boldsymbol{\chi})$ with respect to $\chi_{1}$ and $\chi_{2}$, respectively, and by setting $\chi_{1}=0=\chi_{2}$. The statistics of the work $W_{\tau}$ can then be obtained by using the first law of thermodynamics:
@@ IV.1 The signal-to-noise ratio
Figure 2: Plot of the $\text{SNR}(\Delta U_{\tau})$ from Eq. (6) as a function of the time window $g\tau$, related to four different initial cavity states: pure Fock (continuous green line), coherent (dashed red), thermal (dotted blue), and squeezed coherent (dash-dotted orange). Importantly for a fair comparison between them, all these states have the same average photon number, which is $\langle n\rangle=N=5$. Moreover, here we set the other parameters to be $\omega_{\text{qub}}=1\text{ GHz}$, $g/\omega_{\text{qub}}=10^{-2}$, $\Delta\omega=\omega_{\text{qub}}-\omega_{\text{cav}}$ and $\Delta\omega/\omega_{\text{qub}}=5\cdot 10^{-3}$, $\beta=\log(\frac{N+1}{N})/\hbar\omega_{\text{cav}}$, $\alpha=\sqrt{N}$, $\zeta=0.6$, and $\tilde{\alpha}\simeq 3.906$ (in order to maintain the constraint $\langle n\rangle=N=5$). As reported in Eq. (13), $\text{SNR}(\Delta U_{\tau})$, $\text{SNR}(Q_{\tau})$, and $\text{SNR}(W_{\tau})$ are quantitatively identical.
@@
Analogously, we can calculate the SNR for different initial cavity states, e.g., the ones defined in Eq. (2). Although an analytical formulation is still possible, in these cases it involves summations over infinite terms (i.e., over the cavity energy levels): therefore, computations must be carried out numerically, by truncating the summations. The results corresponding to four different initial cavity mode states are reported in Fig. 2. Here we have fixed $N=5$, and consequently we choose $\beta=\log(\frac{N+1}{N})/\hbar\omega_{\text{cav}}$ (by direct application of the Bose-Einstein statistical relation), and $\alpha=\sqrt{N}$; regarding the squeezed coherent state, we choose a real squeezing parameter $\zeta=0.6$, and we picked $\tilde{\alpha}\simeq 3.906$. In this way, each Gaussian state contains $\langle n\rangle=N=5$ photons on average. Therefore, a direct comparison can be made between the different cases, showing that an effective quantum advantage in the charging process of a JC quantum battery is achieved when the cavity mode is prepared in a Fock state as compared to any Gaussian (i.e., thermal, coherent, or squeezed coherent) states of light, as anticipated at the beginning of this Section.
@@ IV.2 Charging protocol: power and charge efficacy
The simulated results showing the produced power and the fidelity evaluated over a time window $g\tau$ are depicted in Fig. 3. First of all, we immediately notice how the Fock state cavity produces slightly more power than the three Gaussian states. (b) Fidelity $F$, as defined in Eq. (17), between $\hat{\rho}_{\text{qub}}(\tau)$ and $\hat{\rho}_{\text{qub}}^{\text{exc}}$, where $\hat{\rho}_{\text{qub}}^{\text{exc}}=\ket{e}\bra{e}$ |. The optimal charge (i.e., the process for which the final qubit state approaches |$\ket{e}\bra{e}$ |), can be obtained only when the cavity is prepared in a Fock state. In that case, $F\to 1$ when the qubit and the cavity mode are perfectly resonant, and when the charging time is the optimal time of Eq. (15). It is worth emphasising that the charging efficacy is maximized at the very same value of time $\tau_{\text{opt}}$ that optimized the SNR. This fact, far from obvious, strongly stresses the importance of incorporating the role of energy fluctuations into the picture when characterizing the performance of thermodynamic processes. Finally, since our goal is not to characterize the thermodynamics of a generic process, but rather to focus on the specific charging protocol which brings any qubit initial state to the target ‘charged’ state |$\ket{e}\bra{e}$ |, the injected energy $\langle\Delta U\rangle$ can be seen to correspond to the more commonly considered ergotropy [54, 23, 22, 55] (i.e., the maximum extractable work from a system via a unitary operation).
@@ V Discussion
It is worth remarking that, within the quantum batteries community, the expression “advantage” usually refers to a collective enhancement of the charging power (also defined the superlinear scaling of power [56, 57, 58]), which is an intrinsically many-body effect. Instead, the advantage that we investigated is primarily referred to a single-body quantum battery (a single qubit), and most of all to the efficacy of a single charging process. The multi-body advantage is currently under investigation.
Finally, we would like to emphasize that our model could be employed to describe physical setups involving flying qubits (e.g., atoms or ions flying through the confined electromagnetic field mode of an open resonator) as well as static ones (for example, superconducting qubits in circuit QED, or trapped charged ions with sufficiently long trapping times, or even semiconductor quantum dots in high-quality microresonators). In this respect, we believe that an implementation based on flying qubits would be more suitable for the realization of a valid quantum battery, since it would be possible to charge multiple qubits without resetting them immediately after the process. The reset would be unavoidable, instead, in the case of a static qubit: it has to be brought back to the ground state after each iteration step, in order to be recycled to perform a multiple charging process. Nevertheless, a physical setup based on static qubits could still be exploited to obtain a proof-of-principle demonstration of the proposed protocol.
@@ VI Conclusions
As a final consideration, we reckon that these findings can be particularly suitable for the thermodynamic analysis of a wide variety of quantum processes, especially the ones for which energy fluctuations play a prominent role, e.g. in quantum catalytic processes [65]. Moreover, they provide a contribution to recent results concerning the investigation of the energetics of bipartite systems [66, 67], which fit well with the model discussed in the present work.
@@ Acknowledgements
The authors are grateful to S. P. Prasad, M. Maffei, A. Auffèves, M. Kolář, M. B. Arjmandi, L. Slodička, G. Benenti, and L. Razzoli for several fruitful scientific discussions. D.R. acknowledges the kind hospitality of the Majulab at the Centre for Quantum Technologies (CQT) in Singapore, as well as the warm welcome of the Optics Department at the Palacký University in Olomouc. D.G. acknowledges the ‘National Quantum Science Technology Institute’ (NQSTI, PE4, Spoke 1) within the PNRR project PE0000023. G.G. acknowledges funding from the Italian Ministry of Research (MUR) under the “Rita Levi Montalcini” project THESEUS. R.F. acknowledges the grant 23-06308S from the Czech Science Foundation, Project CZ.02.01.01/00/22-008/0004649 of the MEYS Czech Republic supported by the EU funding. R.F. also acknowledges funding from the MEYS of the Czech Republic (Grant Agreement 8C22001), and Project SPARQL funded from the EU Horizon 2020 Research and Innovation Programme under Grant Agreements No. 731473 and 101017733 (QuantERA).
@@ Appendix A The JC evolution operator
$\begin{split}\hat{\mathcal{U}}_{00}&=e^{-\frac{i}{2}\omega_{\text{cav}}(\hat{\mathds{I}}+2\hat{a}^{\dagger}\hat{a})\tau}\biggl{[}\mathcal{C}(\tau,\hat{\varphi}+g^{2}\hat{\mathds{I}})-i\frac{\Delta\omega}{2}\mathcal{S}(\tau,\hat{\varphi}+g^{2}\hat{\mathds{I}})\biggr{]}\\ \hat{\mathcal{U}}_{01}&=-ige^{-\frac{i}{2}\omega_{\text{cav}}\tau}e^{-i\omega_{\text{cav}}\hat{a}^{\dagger}\hat{a}\tau}\mathcal{S}(\tau,\hat{\varphi}+g^{2}\hat{\mathds{I}})\hat{a}\\ \hat{\mathcal{U}}_{10}&=-ige^{+\frac{i}{2}\omega_{\text{cav}}\tau}e^{-i\omega_{\text{cav}}\hat{a}^{\dagger}\hat{a}\tau}\mathcal{S}(\tau,\hat{\varphi})\hat{a}^{\dagger}\\ \hat{\mathcal{U}}_{11}&=e^{+\frac{i}{2}\omega_{\text{cav}}\tau}e^{-i\omega_{\text{cav}}\hat{a}^{\dagger}\hat{a}\tau}\biggl{[}\mathcal{C}(\tau,\hat{\varphi})+i\frac{\Delta\omega}{2}\mathcal{S}(\tau,\hat{\varphi})\biggr{]}\end{split}$ (18)
@@ Appendix B The generating function of the statistics
$\begin{split}&\mathcal{G}_{\tau}(\boldsymbol{\chi})=\rho_{00}(0)\langle\zeta(\hat{a}^{\dagger}\hat{a}+\hat{\mathds{I}})\rangle_{\text{cav}}\\ &+ig\rho_{01}(0)e^{+i\frac{\chi_{1}}{2}\hbar\omega_{\text{qub}}+i\frac{\chi_{2}}{2}\hbar\omega_{\text{cav}}}\langle\hat{a}^{\dagger}\eta(\hat{a}^{\dagger}\hat{a}+\hat{\mathds{I}})\rangle_{\text{cav}}\\ &-ig\rho_{10}(0)e^{+i\frac{\chi_{1}}{2}\hbar\omega_{\text{qub}}+i\frac{\chi_{2}}{2}\hbar\omega_{\text{cav}}}\langle\eta^{\ast}(\hat{a}^{\dagger}\hat{a}+\hat{\mathds{I}})\hat{a}\rangle_{\text{cav}}\\ &+g^{2}\rho_{11}(0)e^{+i\chi_{1}\hbar\omega_{\text{qub}}+i\chi_{2}\hbar\omega_{\text{cav}}}\langle\hat{a}^{\dagger}\mathscr{S}^{2}(\tau,\hat{a}^{\dagger}\hat{a}+\hat{\mathds{I}})\hat{a}\rangle_{\text{cav}}\\ &+g^{2}\rho_{00}(0)e^{-i\chi_{1}\hbar\omega_{\text{qub}}-i\chi_{2}\hbar\omega_{\text{cav}}}\langle\hat{a}\mathscr{S}^{2}(\tau,\hat{a}^{\dagger}\hat{a})\hat{a}^{\dagger}\rangle_{\text{cav}}\\ &-ig\rho_{01}(0)e^{-i\frac{\chi_{1}}{2}\hbar\omega_{\text{qub}}-i\frac{\chi_{2}}{2}\hbar\omega_{\text{cav}}}\langle\eta(\hat{a}^{\dagger}\hat{a})\hat{a}^{\dagger}\rangle_{\text{cav}}\\ &+ig\rho_{10}(0)e^{-i\frac{\chi_{1}}{2}\hbar\omega_{\text{qub}}-i\frac{\chi_{2}}{2}\hbar\omega_{\text{cav}}}\langle\hat{a}\eta^{\ast}(\hat{a}^{\dagger}\hat{a})\rangle_{\text{cav}}\\ &+\rho_{11}(0)\langle\zeta(\hat{a}^{\dagger}\hat{a})\rangle_{\text{cav}}\end{split}$ (23)
@@
and $\langle\hat{A}\rangle_{\text{cav}}\equiv\text{Tr}_{\text{cav}}\bigl{\{}\hat{A}(\tau)\hat{\rho}_{\text{cav}}(0)\bigr{\}}$. In our notation, $f(\hat{a}^{\dagger}\hat{a}+\hat{\mathds{I}})$ means that the substitution $\hat{\varphi}\mapsto\hat{\varphi}+g^{2}\hat{\mathds{I}}$ must be performed into definitions (24).
Expression (23) greatly simplifies if we choose, for example, $\hat{\rho}_{\text{qub}}(0)=\ket{g}\bra{g}$ |: by adopting the convention |$\ket{e}=\footnotesize\begin{pmatrix}1\\ 0\end{pmatrix}$ and |$\ket{g}=\footnotesize\begin{pmatrix}0\\ 1\end{pmatrix}$, indeed, $\rho_{ij}(0)=1\iff i=j=1$, otherwise $\rho_{ij}(0)=0$ for $i\neq j$ or $i=j=0$. This is actually the case of the charging protocol of the JC quantum battery, where the qubit starts in the ground state. Now, let we further assume that the initial state of the cavity is a Fock state $\hat{\rho}_{\text{cav}}(0)=\ket{N}\bra{N}$ | with $N$ photons, and let us set $\chi_{2}=0$ (for example, because we may be interested in computing the statistics of $\Delta U_{\tau}$ only). In that case, we get Eq. (7). This can be easily seen from Eq. (23), which becomes
$\begin{split}\mathcal{G}_{\tau}(\boldsymbol{\chi})&=g^{2}e^{+i\chi_{1}\hbar\omega_{\text{qub}}}\langle\hat{a}^{\dagger}\mathscr{S}^{2}(\tau,\hat{a}^{\dagger}\hat{a})\hat{a}\rangle_{\text{cav}}\\ &\quad+\langle\zeta(\hat{a}^{\dagger}\hat{a})\rangle_{\text{cav}}\end{split}$ (25)
because the only non-vanishing factor is $\rho_{11}(0)$. The operator-valued expression $\zeta(\hat{a}^{\dagger}\hat{a})$, defined in Eq. (24), as well as the expression $\hat{a}^{\dagger}\mathscr{S}^{2}(\tau,\hat{a}^{\dagger}\hat{a})\hat{a}$, are then traced over the cavity’s degrees of freedom, together with the initial state of the cavity. For what concerns $\zeta(\hat{a}^{\dagger}\hat{a})$, we have $\langle\zeta(\hat{a}^{\dagger}\hat{a})\rangle_{\text{cav}}=\text{Tr}\{\zeta(\hat{a}^{\dagger}\hat{a})\ket{N}\bra{N}\}$ | }, that translates into $\bra{N}\zeta(\hat{a}^{\dagger}\hat{a})\ket{N}$. It is then straightforward to evaluate $\bra{N}\mathcal{S}^{2}(\tau,\hat{\varphi})\ket{N}=\mathscr{S}^{2}(\tau)$ and $\bra{N}\mathcal{C}^{2}(\tau,\hat{\varphi})\ket{N}=\mathscr{C}^{2}(\tau)$, where the functions $\mathscr{S}(\tau)$ and $\mathscr{C}(\tau)$ are respectively defined as: